OpenStudy (anonymous):

find the domain, points of discontinuity, and x and y intercepts of each rational function. determine whether the discontinuities are removable or non removable. y= x^2 + 2x/ x^2 +2

OpenStudy (turingtest):

\[y={x^2+2x\over x^2+2}\]?

OpenStudy (anonymous):

yes, sorry i don't know how to do that line thing

OpenStudy (turingtest):

it's okay, just next time use parentheses y=(x^2+2x)/(x^2+2) would have been much clearer anyway, first is domain: are there any values for which this function is undefined?

OpenStudy (anonymous):

i don't think so.

OpenStudy (turingtest):

right, the denominator is never zero, we have no negative square roots or logarithms so it is defined for all reals...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and i think the points of discontinuity are x=1, and x =6

OpenStudy (anonymous):

?

OpenStudy (turingtest):

why?

OpenStudy (anonymous):

oh no sorry. i was looking at wrong problem

OpenStudy (turingtest):

discontinuity: it's defined for all reals, and has no "jump" points, so it is continuous everywhere

OpenStudy (turingtest):

do you know how to find x and y intercepts?

OpenStudy (anonymous):

yeah don't you set x= to 0 and y = to 0 and substitute it into the equation ?

OpenStudy (turingtest):

exactly

OpenStudy (anonymous):

so there are no points of discontinuity then?

OpenStudy (turingtest):

no if the function is defined everywhere, and is not a step function, then it is (as far as I know) always continuous everywhere

OpenStudy (anonymous):

okay

OpenStudy (turingtest):

that makes the third question trivial, since we have no discontinuities to talk about so I guess we're done :)

OpenStudy (turingtest):

or fifth Q or whatever it was

OpenStudy (anonymous):

okay. thanks.

OpenStudy (turingtest):

anytime :)

OpenStudy (anonymous):

hey would you kknow how to factor x^2+2 or x^2 + 9 i forgot?

OpenStudy (turingtest):

neither can be factored by "traditional" means x^2-9=(x+3)(x-3) ^^^^^^^^^perhaps that is what you are thinking of?

OpenStudy (anonymous):

find the vertical asymptotes and holes for the graph of each rational function. y=(x+5) / (x^2+9) and when i do these types of problems usually i factor it out to get the asymptotes but idk how to factor the bottom. please help?